### v3

Abstract
In 1983, Hartle and Hawking proposed the no-boundary proposal, suggesting that the universe has no beginning in the sense of a spacetime singularity or boundary. Nevertheless, there is an origin of time. Mathematically, this involves signature-type changing manifolds in which a Riemannian region smoothly transitions to a Lorentzian region across the hypersurface $\mathcal{H}$ where time begins.
We develop a coherent framework for signature changing manifolds with a degenerate yet smooth metric. Established Lorentzian tools and results are then adapted to this setting, and new definitions are introduced that carry unforeseen causal implications. A noteworthy consequence is the presence of locally time-reversing loops through every point on the hypersurface. Imposing global hyperbolicity on the Lorentzian region, we prove that for every point $p\in M$ there exists a pseudo-timelike loop self-intersecting at $p$. Equivalently, $M$ always admits a closed pseudo-timelike path around which the time direction reverses, preventing any consistent distinction between future- and past-directed vectors. To an observer near $\mathcal{H}$, such loops may appear as the creation of a particle-antiparticle pair at two distinct points.
keywords:
causality, pseudo-timelike curve, singular semi-Riemannian geometry, Lorentzian geometry, signature change, mathematical general relativity, singular metric
1 Introduction
According to popular ideas about quantum cosmology, classical cosmological models contain an initial Riemannian region of Euclidean signature joined to a final semi-Riemannian region with the usual Lorentzian signature [12, 13]. In 1983 Hartle and Hawking [22] introduced a conceptually intriguing idea involving signature-type change, which led to the no-boundary proposal for the initial conditions of the universe. [^1] Although the Hartle-Hawking proposal does not explicitly address signature change, it naturally motivates the concept: the universe “tunnels” into existence from a state with no time, a Euclidean-signature spacetime without a past boundary, to a Lorentzian spacetime. This process effectively eliminates the classical Big Bang singularity. The resulting transition and its classical implications have been discussed in [12, 36]. In this proposal, the universe has no initial singularity or boundary. [^2] Although singularities can be considered points where curves terminate at finite parameter values, providing a general definition remains difficult [14]. While such a spacetime is singularity-free and thus has no distinct beginning, it still possesses an origin of time [12].
Since a signature-type changing metric is necessarily either degenerate or discontinuous at the locus of signature change [9], we will allow for the metric to become degenerate. Hence, in the present article we will discuss singular semi-Riemannian manifolds for which the metric constitutes a smooth $(0,2)$-tensor field that is degenerate at a subset $\mathcal{H}\subset M$, where the bilinear type of the metric changes upon crossing $\mathcal{H}$.
Although the compatibility of the Riemannian and Lorentzian domains is assumed to be established, insofar as the metric should be smooth on the interface $\mathcal{H}$, the behavior of curves as they cross this interface still requires further study. Moreover, in a manifold where the signature changes from (+,+,,+)($\ldots$,+) to (,+,,+)(-$\ldots,+$), the conventional concept of timelike (or spacelike) curves does not exist anymore. This gives rise to a new notion of curves called pseudo-timelike and pseudo-spacelike curves. In order to define these curves we make a detour to draw upon the concept of the generalized affine parameter which we use as a tool to distinguish genuine pseudo-timelike (and pseudo-spacelike, respectively) curves from curves that asymptotically become lightlike as they approach the hypersurface of signature change.
We endeavor to adapt well-established Lorentzian tools and results to the signature changing setting, as far as possible. This task proves to be less straightforward than anticipated, necessitating the introduction of new definitions with unexpected causal implications, reaching a critical juncture in our exploration. We draw upon the definition of pseudo-time orientability and the given absolute time function to decide whether a pseudo-timelike curve is future-directed. This establishes the definition for the pseudo-chronological past (and pseudo-chronological future) of an event.
In this article we show that for signature-type change of the delineated type, all these considerations lead to a surprising theorem revealing the non-well-behaved nature of these manifolds. In a sufficiently small region near the junction of signature change $\mathcal{H}$, transverse signature-type changing manifolds with a transverse radical exhibit local anomalies: Specifically, each point on the junction facilitates a closed time-reversing loop, challenging conventional notions of temporal consistency. [^3] In more informal terms, in general relativity, a closed timelike curve is a smooth, timelike loop where, at every intersection point, the direction of movement is consistently the same. In contrast, a loop is a broader concept where a timelike curve loops back on itself, but the direction of movement at the intersection points might not always be the same. This is a more intuitive explanation; for a precise mathematical definition and its extension to a setting with signature-type change, see Definition 9. Or put differently, there always exists a closed pseudo-timelike path in $M$ along which the direction of time reverses, making it impossible to consistently assign future- and past-directed vectors. By imposing global hyperbolicity on the Lorentzian region, the global analog can be proven by showing that for every point $p\in M$, there exists a pseudo-timelike loop intersecting at $p$. In other words, a transverse, signature-type changing manifold with a transverse radical always admits pseudo-timelike loops. From the viewpoint of an observer in the Lorentzian region near the hypersurface $\mathcal{H}$, such locally pseudo-timelike loops may be interpreted as the creation of a particle-antiparticle pair at two distinct points $\hat{q},q\in\mathcal{H}$.
1.1 Transverse type-changing singular semi-Riemannian manifold
Unless otherwise specified, the considered manifolds, denoted as $M$ with dimension $\dim(M)=n$, are assumed to be locally homeomorphic to $\mathbb{R}^{n}$. Moreover, these manifolds are expected to be connected, second countable, and Hausdorff. This definition also indicates that all manifolds have no boundary. Additionally, we will generally assume that the manifolds under consideration are smooth. Unless stated otherwise, all related structures and geometric objects (such as curves, maps, fields, differential forms, etc.) are assumed to be smooth as well.
Definition 1.
A singular semi-Riemannian manifold is a generalization of a semi-Riemannian manifold. It is a differentiable manifold having on its tangent bundle a symmetric bilinear form which is allowed to become degenerate.
Definition 2.
Let $(M,g)$ be a singular semi-Riemannian manifold and let be $p\in M$. We say that the metric changes its signature at a point $p\in M$ if any neighborhood of $p$ contains at least one point $q$ where the metric’s signature differs from that at $p$.
We align with [25] in requiring that $(M,g)$ be a semi-Riemannian manifold with $\dim M\geq 2$, where $g$ is a smooth, symmetric, degenerate $(0,2)$-tensor on $M$, and $\mathcal{H}:=\{q\in M\!\!:g\!\!\mid_{q}is\;degenerate\}$. This means $\mathcal{H}$ is the locus where the rank of $g$ fails to be maximal. In addition, we assume that one connected component of $M\setminus\mathcal{H}$ is Riemannian, denoted by $M_{R}$, while all other connected components $(M_{L_{\alpha}})_{\alpha\in I}\subseteq M_{L}\subset M$ are Lorentzian, where $M_{L}:=\underset{\alpha\in I}{\bigcup}M_{L_{\alpha}}$represents the Lorentzian domain. Furthermore, we assume throughout that the point set $\mathcal{H}$, where $g$ becomes degenerate is not empty.
Remark 1.
This restriction permits transitions of the form $M_{L}-M_{R}-M_{L}$; however, it forbids $M_{R}-M_{L}-M_{R}$, which could correspond to transitions involving both an initial and a final singularity. This limitation is of a physical nature, ensuring that the Hartle-Hawking “no boundary" proposal is satisfied, thereby allowing the Lorentzian part $M_{L}$ to be interpreted as our universe. Then within $M_{L}$, Einstein’s equations hold—these form a hyperbolic system of partial differential equations whose solution (given known matter conditions, such as a vacuum) is determined by a Riemannian hypersurface and the time derivatives of the induced metric on that hypersurface.
On the other hand, hypersurfaces that permit a signature change are highly constrained, as the evolution of the solution for $g$ would have to degenerate across an entire hypersurface. Consequently, we assume that the Lorentzian part cannot revert to a region of Riemannian nature. For the Riemannian part, this restriction does not apply. First, there is no reason to assume Einstein’s equations for conventional matter in this region. Second, the system would no longer be hyperbolic in nature. From a purely mathematical perspective, this argument is, of course, irrelevant—at the level of the metric $g$, multiple transitions can certainly be allowed.
Moreover, we impose the following two conditions [25]:
  1. 1.
    We call the metric $g$ a codimension-$1$ transverse type-changing metric if $d(\det([g_{\mu\nu}]))_{q}\neq 0$ for any $q\in\mathcal{H}$ and any local coordinate system $\xi=(x^{0},\ldots,x^{n-1})$ around $q$. Then we call $(M,g)$ a transverse type-changing singular semi-Riemannian manifold [1, 2, 25]
    This implies that the subset $\mathcal{H}\subset M$ is a smoothly embedded hypersurface in $M$, and the bilinear type of $g$ changes upon crossing $\mathcal{H}$. Moreover, at every point $q\in\mathcal{H}$ there exists a one-dimensional subspace, denoted as the radical $Rad_{q}\subset T_{q}M$, within the tangent space $T_{q}M$ that is orthogonal to all of $T_{q}M$ at that point.
  2. 2.
    The radical $Rad_{q}$ is transverse to $\mathcal{H}$ for any $q\in\mathcal{H}$. In cosmological applications, particularly in the context of “no boundary” models, a spacelike surface of signature change is the natural and preferred choice. Thus, our focus is specifically directed toward the concept of a transverse radical, which ensures such a spacelike hypersurface of signature change. Henceforward, we assume throughout that $(M,g)$ is a singular transverse type-changing semi-Riemannian manifold with a transverse radical, unless explicitly stated otherwise. [^4] In the case of a tangential radical, no statements corresponding to the main results of our work can be made, since in $M_{L}$ causal curves approaching the hypersurface tend to become tangent to it rather than intersecting it transversely. This corresponds precisely to the case of a lightlike surface, as in some models of signature change in black hole spacetimes.
Remark 2.
Recall that the radical at $q\in\mathcal{H}$ is defined as the subspace $Rad_{q}:=\{w\in T_{q}M:g(w,\centerdot)=0\}$. This means $g(v_{q},\centerdot)=0$ for all $v_{q}\in Rad_{q}$. Note that the radical can be either transverse or tangent to the hypersurface $\mathcal{H}$. The radical $Rad_{q}$ is called transverse [26] if $Rad_{q}$ and $T_{q}\mathcal{H}$ span $T_{q}M$ for any $q\in\mathcal{H}$, i.e. $Rad_{q}+T_{q}\mathcal{H}=T_{q}M$. This means that $Rad_{q}$ is not a subset of $T_{q}\mathcal{H}$, and obviously, $Rad_{q}$ is not tangent to $\mathcal{H}$ for any $q$.
The following theorem is a direct consequence of the above two conditions and the results in [25, 27], but it is restated here in a more concise and self-contained form for clarity. Moreover, we have refined its statement for improved readability and conceptual transparency:
Theorem 1.
Let $M$ be a singular semi-Riemannian manifold endowed with a $(0,2)$-tensor field $g$ and the surface of signature change defined as $\mathcal{H}:=\{q\in M\!\!:g\!\!\mid_{q}is\;degenerate\}$. Then $(M,g)$ is a transverse, signature-type changing manifold with a transverse radical if and only if for every $q\in\mathcal{H}$ there exist a neighborhood $U(q)$ and smooth coordinates $(t,x^{1},\ldots,x^{n-1})$ such that $g=-t(dt)^{2}+g_{ij}(t,x^{1},\ldots,x^{n-1})dx^{i}dx^{j}$, for $i,j\in\{1,\ldots,n-1\}$.
In the style of time-orthonormal coordinates in Lorentzian geometry we denote the coordinates in Theorem 1 as radical-adapted Gauss-like coordinates. It is now possible to simplify matters by using these coordinates whenever dealing with a transverse, signature-type changing manifold with a transverse radical. Notably, signature-type change and the radical-adapted Gauss-like coordinates imply the existence of an uniquely determined, coordinate independent, natural absolute time function $\mathfrak{h}(t,\mathbf{\hat{x}}):=t$, with $\mathbf{\hat{x}}=(x^{1},\ldots,x^{n-1})$, in the neighborhood of the hypersurface [25]. Then the absolute time function establishes a foliation [17, 21] in a neighborhood of $\mathcal{H}$, such that $\mathcal{H}$ is a level surface of that decomposition.
1.2 Statement of results
Before presenting the main results we require some new definitions.
Definition 3 (Pseudo-timelike curve).
Let $\gamma\colon[a,b]\to M$ be a continuous and differentiable curve, with $a,b]\subset\mathbb{R}$ and $\infty<a<b<\infty$. We call $\gamma=\gamma^{\mu}(u)=x^{\mu}(u)$ in $M$ a pseudo-timelike (respectively, pseudo-spacelike) curve if
  1. 1.
    $\operatorname{Im}(\gamma)\cap M_{L}\neq\varnothing$, i.e. $\gamma$ has image points in the Lorentzian region; and
  2. 2.
    for every generalized affine parametrization (see Definition 6) of $\gamma$ in $M_{L}$ there exists $\varepsilon>0$ such that $g(\gamma^{\prime},\gamma^{\prime})<-\varepsilon$ (respectively, $g(\gamma^{\prime},\gamma^{\prime})>\varepsilon$).
In simpler terms, we call a curve pseudo-timelike if it is timelike in the Lorentzian domain $M_{L}$ and does not become asymptotically lightlike as it approaches the hypersurface where the signature changes. Consequently, a pseudo-timelike loop is a generalization of a pseudo-timelike curve that loops back on itself. However, unlike a regular closed curve where the direction of movement would be the same at every intersection point, in a pseudo-timelike loop, the direction of movement at the intersection points is not necessarily the same (see Definition 9).
Definition (Pseudo-timelike).
A vector field $V$ on a signature-type changing manifold $(M,g)$ is pseudo-timelike if and only if its integral curves are pseudo-timelike (hence, in particluar, V\,V is timelike in $M_{L}$).
Definition (Pseudo-time orientable).
A signature-type changing manifold $(M,g)$ is pseudo-time orientable if and only if the Lorentzian region $M_{L}$ is time orientable.
In a sufficiently small region near the junction of signature change, transverse, signature-type changing manifolds with a transverse radical exhibit local anomalies. Specifically, each point on the junction gives rise to the existence of closed time-reversing loops, challenging conventional notions of temporal consistency.
Theorem 2 (Local loops).
[…]
The existence of such pseudo-timelike curves locally near the hypersurface that loop back to themselves, gives naturally rise to the question whether this type of curves also occur globally. In the global version a key notion is global hyperbolicity which plays a role in the spirit of completeness for Riemannian manifolds. By imposing the constraint of global hyperbolicity on the Lorentzian region, we demonstrate
Theorem 3 (Global loops).
[…]
2 Pseudo-causal and pseudo-lightlike curves
In $(M,g)$ be an $n$-dimensional manifold on which the metric signature changes from Riemannian, (+,+,,+)($\text{$\cdots$}$,+), to Lorentzian, (,+,,+)($\text{\textminus}$$\text{$\cdots$},+$). In such a setting, the conventional notion of timelike curves ceases to apply globally. From a point located in the Lorentzian region, the hypersurface of signature change may be reached within finite proper time. However, in the Riemannian region no meaningful notion of proper time exists, and hence curves there cannot be classified as timelike, spacelike, or null. Consequently, in manifolds with signature change this gives rise to a novel notion of curves. In order to define those curves we have to make a detour to draw upon the concept of the generalized affine parameter.
2.1 Properties of the generalized affine parameter
In this section we introduce the notions of pseudo-timelike and pseudo-spacelike curves. To make these notions well-defined, we must distinguish genuine pseudo-timelike (respectively pseudo-spacelike) curves from those that asymptotically become lightlike as they approach the hypersurface of signature change. For this purpose, the concept of a generalized affine parameter will play a central role. In particular, we require a suitable notion of completeness, ensuring that every $C^{1}$ curve of finite length with respect to such a parameter possesses an endpoint. The use of generalized affine parameters in this context was first proposed by Ehresmann [11] and later developed further by Schmidt [35], who employed them to characterize the completeness of arbitrary curves. The generalized affine parameter proves especially useful quantity for probing singularities, because it can be defined for an arbitrary curve, not necessarily a geodesic.
Definition 6 (Generalized affine parameter).
Let $M$ be an $n$-dimensional manifold with an affine connection, and let $\gamma\colon J\rightarrow M$ be a $C^{1}$ curve. A smooth vector field $V$ along $\gamma$ is a smooth map $V\colon J\rightarrow TM$ such that $V(t)\in T_{\gamma(t)}M$ for all $t\in J$. Such a vector field $V$ along $\gamma$ is called parallel along $\gamma$ if it satisfies the differential equation
$\nabla_{\gamma^{\prime}}V(t)=0\quad\text{for all }t\in J$
(see [3] for details).
Fix $t_{0}\in J$ and and set $p_{0}=\gamma(t_{0})$. Choose a basis $\{e_{1},\dots,e_{n}\}$ of $T_{\gamma(t_{0})}M$. Let $E_{i}$ be the unique parallel field along $\gamma$ with $E_{i}(t_{0})=e_{i}$ for $1\leq i\leq n$. Then $\{E_{1}(t),E_{2}(t),\ldots,E_{n}(t)\}$ forms a basis for $T_{\gamma(t)}M$ for each $t\in J$. The tangent vector of $\gamma$ at $p_{0}$, can be expressed in this frame as a linear combination of the elements of the chosen basis with coefficients
$V^{i}(t):\dot{\gamma}(t)=\sum_{i=1}^{n}\underset{V^{i}(t)E_{i}(t)}{\underbrace{V^{i}(t)E_{\gamma(t)i}}}$,
where $V^{i}\colon J\longrightarrow\mathbb{R}$ for $1\leq i\leq n$. Then the generalized affine parameter $\mu=\mu(\gamma,E_{1},\ldots,E_{n})$ of $\gamma(t)$ associated with this basis is given by
$\mu(t)=\int_{t_{0}}^{t}\sqrt{\mathop{\sum_{i=1}^{n}[V^{i}(t)]^{2}}}dt=\int_{t_{0}}^{t}\sqrt{\delta_{ij}V^{i}(t)V^{j}(t)}dt,\;t\in J.$ (1)
The assumption that $\gamma$ is $C^{1}$ ensures that the parallel fields $\{E_{1},\dots,E_{n}\}$ are well-defined.
Furthermore, we have
Proposition 4.
[24] The curve $\gamma$ has a finite arc-length in the generalized affine parameter $\mu=\mu(\gamma,E_{1},\ldots,E_{n})$ if and only if $\gamma$ has finite arc-length in any other generalized affine parameter $\mu=\mu(\gamma,\bar{E}_{1},\ldots,\bar{E}_{n})$.
Note that the generalized affine parameter of a curve depends on the chosen basis. Conceptually, one treats the parallel-transported basis vectors as if they were orthonormal with respect to a Riemannian metric, and then defines the “length” of $\gamma(t)$ accordingly. In particular, if the metric $g$ is positive definite, the generalized affine parameter associated with an orthonormal basis coincides with the usual arc-length.
This notion of completeness succeeds in distinguishing precisely the behavior we intend to capture. A further advantage is that $\mu$ can be defined for any $C^{1}$ curve, including null curves as well as timelike or spacelike ones. Moreover, whenever a curve has unbounded proper length, its generalized affine parameter is necessarily unbounded as well [3].
Claim 1.
If the generalized affine parameter of a curve is finite for one choice of parallel frame, then it is finite for every such choice. Hence, for completeness it suffices to know whether the generalized affine parameter is finite or infinite, regardless of the chosen frame.
Proof.
Let $\{E_{i}\}$ and $\{\tilde{E}_{j}\}$ be two parallel bases of $T_{\gamma(t)}M$ along $\gamma$, and the components $V^{i}(t)$ with respect to another basis are given by $\tilde{V}^{j}(t)=\sum_{i=1}^{n}A_{i}^{j}V^{i}(t)$. Then
$\dot{\gamma}(t)=\sum_{i=1}^{n}V^{i}(t)E_{i}(t)=\sum_{j=1}^{n}\tilde{V}^{j}(t)\tilde{E}_{j}(t),$
with $\tilde{V}^{j}(t)=\sum_{i=1}^{n}A_{i}^{j}V^{i}(t)$, where $A=(A_{i}^{j})$ is a constant invertible matrix. Hence also $V^{j}(t)=\sum_{i=1}^{n}a_{i}^{j}\tilde{V}^{i}(t)$ with $A^{-1}=(a_{i}^{j})$. Then the generalized affine parameters with respect to these basis are $\mu(t)=\int_{t_{0}}^{t}\sqrt{\mathop{\sum_{i=1}^{n}[V^{i}(t)]}^{2}}dt$ and $\tilde{\mu}(t)=\int_{t_{0}}^{t}\sqrt{\mathop{\sum_{i=1}^{n}[\tilde{V}^{i}(t)]^{2}}}dt$. By direct estimate,
$\left|\tilde{V}(t)\right|=\left|\sum_{i=1}^{n}A_{i}^{j}V^{i}(t)\right|\leq\sum_{i=1}^{n}\mid A_{i}^{j}\mid\mid V^{i}(t)\mid\leq\underset{ij}{\max}\sideset{\mid A_{i}^{j}\mid}{{}_{i=1}^{n}}{\sum}\mid V^{i}(t)\mid.$
Applying Cauchy–Schwarz gives
$|\tilde{V}^{j}(t)|^{2}\leq\max_{i,j}|A_{i}^{j}|^{2}\underbrace{\left(\sum_{i=1}^{n}|V^{i}(t)|\right)^{2}}_{\left(\sum_{i=1}^{n}|V^{i}(t)|\cdot 1\right)^{2}}\\ \leq\max_{i,j}|A_{i}^{j}|^{2}\left(\sum_{i=1}^{n}|V^{i}(t)|^{2}\right)\left(\sum_{i=1}^{n}1\right)=n\cdot\max_{i,j}|A_{i}^{j}|^{2}\sum_{i=1}^{n}|V^{i}(t)|^{2}.$
Summing over $j$ yields
$\sum_{j=1}^{n}|\tilde{V}^{j}(t)|^{2}\;\leq\;\sum_{j=1}^{n}\Biggl(n\cdot\max_{i,j}|A_{i}^{j}|^{2}\sum_{i=1}^{n}|V^{i}(t)|^{2}\Biggr)\;=\;n^{2}\cdot\max_{i,j}|A_{i}^{j}|^{2}\sum_{i=1}^{n}|V^{i}(t)|^{2}.$
On the other hand, we get $\sum_{j=1}^{n}|V^{j}(t)|^{2}\;\leq\;n^{2}\max_{i,j}|a_{i}^{j}|^{2}\sum_{i=1}^{n}|\tilde{V}^{i}(t)|^{2}.$ Combining both estimates yields
$\sum_{j=1}^{n}|\tilde{V}^{j}(t)|^{2}\leq n^{2}\cdot\max_{i,j}|A_{i}^{j}|^{2}\left(\sum_{i=1}^{n}|V^{i}(t)|^{2}\right)\leq n^{4}\cdot\max_{i,j}|A_{i}^{j}|^{2}\cdot\max_{i,j}|a_{i}^{j}|^{2}\left(\sum_{i=1}^{n}|\tilde{V}^{i}(t)|^{2}\right).$
This is equivalent to
$\frac{1}{n^{2}\max_{i,j}|A_{i}^{j}|^{2}}\sum_{j=1}^{n}|\tilde{V}^{j}(t)|^{2}\;\leq\;\sum_{i=1}^{n}|V^{i}(t)|^{2}\;\leq\;n^{2}\max_{i,j}|a_{i}^{j}|^{2}\sum_{i=1}^{n}|\tilde{V}^{i}(t)|^{2},$
which implies
$\underbrace{\frac{1}{\sqrt{\,n^{2}\max_{i,j}|A_{i}^{j}|^{2}}}}_{c_{1}}\sqrt{\sum_{j=1}^{n}|\tilde{V}^{j}(t)|^{2}}\;\leq\;\sqrt{\sum_{i=1}^{n}|V^{i}(t)|^{2}}\;\leq\;\underbrace{\sqrt{\,n^{2}\max_{i,j}|a_{i}^{j}|^{2}}}_{c_{2}}\sqrt{\sum_{i=1}^{n}|\tilde{V}^{i}(t)|^{2}}.$
Integrating, the generalized affine parameters
$\mu(t)=\int_{t_{0}}^{t}\sqrt{\sum_{i=1}^{n}|V^{i}(s)|^{2}}\,ds,\quad\tilde{\mu}(t)=\int_{t_{0}}^{t}\sqrt{\sum_{j=1}^{n}|\tilde{V}^{j}(s)|^{2}}\,ds$
satisfy
$c_{1}\tilde{\mu}(t)\leq\mu(t)\leq c_{2}\tilde{\mu}(t),$ (2)
for positive constants $c_{1},c_{2}$. Thus, $\mu(t)$ is finite if and only if $\tilde{\mu}(t)$ is finite.
2.2 Application of the generalized affine parameter in a signature-type changing manifold
Let $M=M_{L}\cup\mathcal{H}\cup M_{R}$ be an $n$-dimensional transverse type-changing singular semi-Riemannian manifold with a type-changing metric $g$, and $\mathcal{H}:=\{q\in M\!\!:g\!\!\mid_{q}is\;degenerate\}$ the locus of signature change. We further assume that one component, $M_{L}$, of $M\setminus\mathcal{H}$ is Lorentzian and the other one, $M_{R}$, is Riemannian.
Definition 7 (Pseudo-lightlike curve).
Given a continuous and differentiable curve $\gamma\colon[a,b]\longrightarrow M$, with $a,b]\subset\mathbb{R}$, where $\infty<a<b<\infty$. Then the curve $\gamma=\gamma^{\mu}(u)=x^{\mu}(u)$ is a pseudo-lightlike curve if
  • its tangent vector field in the Lorentzian component $M_{L}$ is null,
  • its tangent vector field in the Riemannian component $M_{R}$ is arbitrary.
A similar definition applies for a pseudo-causal curve. Note that an analogous definition for pseudo-timelike and pseudo-spacelike curves turns out to be problematic as the definition would also include curves that asymptotically become lightlike as they approach $\mathcal{H}$, see Figure 1.
Figure 1: The curve $\gamma$ is not pseudo-timelike since it approaches a null vector at the locus of signature change. This curve is asymptotically lightlike.
Example 1.
Consider the metric $g=t\,(dt)^{2}+(dx)^{2}$ on $\mathbb{R}^{2}$, and the non-parametrized, non-geodesic curve $\gamma$ defined by
$\tan x=\frac{2}{3}\,\mathrm{sgn}(t)\,|t|^{3/2},\quad-\frac{\pi}{2}<x<\frac{\pi}{2}.$
We can rearrange this equation to isolate $x$ (see Figure 2):
$\frac{3}{2}\tan x=\operatorname{sgn}(x)\,\biggl|\frac{3}{2}\tan x\biggr|=\operatorname{sgn}(t)\,|t|^{\frac{3}{2}}$
$\Longleftrightarrow\underbrace{\operatorname{sgn}(t)\,|t|}_{t}=\operatorname{sgn}(x)\,\Bigl|\frac{3}{2}\tan x\Bigr|^{\frac{2}{3}}=t.$
Next, [^5] Only the coordinate transformation for $T$ is taken from [10]. The example itself, including the choice of the curve $\gamma$ and the analysis of $\gamma$, are our original contributions. we reintroduce the coordinate transformation suggested in [10], expressed as
$T=\int_{0}^{t}\sqrt{\mid\tilde{t}\mid}d\tilde{t}=\frac{2}{3}\sqrt{\mid t\mid}^{3}\cdot\textrm{sgn}(t).$
This yields the metric $g=\textrm{sgn}(T)(dT)^{2}+(dx)^{2}$, and for the curve $\gamma$ we have $T=\tan x$. Thus, in the $(T,x)$-coordinates, $\gamma$ is simply the $\tan$-function, with derivative $dT/dx=1/\cos^{2}x$. Consequently, $\gamma$ is timelike in $M_{L}$ , approaching the light cone from timelike infinity and touching it tangentially as $T\rightarrow 0$ (where the derivative becomes $\frac{1}{\cos^{2}(0)}=1$ in the limit). These are precisely the types of curves we aim to exclude in our definition. Note that unlike $(T,x)$, the $(t,x)$-coordinates cover the entire manifold $M$.
Figure 2: The curve defined by $t=\textrm{sgn}(x)\cdot\left(\mid\frac{3}{2}\tan x\mid\right)^{\frac{2}{3}}$.
Moreover, if the curve $\gamma=(T(s),x(s))$ is parametrized by arc length $s$, then in the $(t,x)$-coordinate system both $\frac{dx}{ds}$ and $\frac{dt}{ds}$ diverge in $M_{L}$, since $1=-(\frac{dT}{ds})^{2}+(\frac{dx}{ds})^{2}$:
$1=-\left(\frac{dT}{dx}\frac{dx}{ds}\right)^{2}+\left(\frac{dx}{ds}\right)^{2}=\left(-\left(\frac{d\tan x}{dx}\right)^{2}+1\right)\left(\frac{dx}{ds}\right)^{2}=\left(-\frac{1}{\cos^{4}x}+1\right)\left(\frac{dx}{ds}\right)^{2},$
which is equivalent to $(\frac{dx}{ds})^{2}=\frac{-1}{(-\frac{1}{\cos^{4}x}+1)}$. And this implies
$\underset{x\rightarrow 0}{\lim}\frac{dx}{ds}=\underset{x\rightarrow 0}{\lim}\pm\sqrt{\frac{-1}{(-\frac{1}{\cos^{4}x}+1)}}=\pm\infty.$
Similarly, we get
$-1$ $=\left(-1+\frac{1}{\left(\frac{dT}{dx}\right)^{2}}\right)\left(\frac{dT}{ds}\right)^{2}$
$=\left(-1+\underset{\frac{1}{(\frac{d\tan x}{dx})^{2}}}{\underbrace{\cos^{4}x}}\right)\left(\frac{dT}{dt}\right)^{2}\left(\frac{dt}{ds}\right)^{2}$
$=\left(-1+\frac{1}{(1+\tan^{2}x)^{2}}\right)|t|\left(\frac{dt}{ds}\right)^{2}$
$=\left(-1+\frac{1}{(1+T^{2})^{2}}\right)|t|\left(\frac{dt}{ds}\right)^{2}$
$=\left(-1+\frac{1}{(1+\frac{4}{9}|t|^{3})^{2}}\right)|t|\left(\frac{dt}{ds}\right)^{2}.$ (3)
This is equivalent to
$\left(\frac{dt}{ds}\right)^{2}=\frac{-1}{\left(-1+\frac{1}{\left(1+\frac{4}{9}|t|^{3}\right)^{2}}\right)|t|},$ (4)
and then yields
$\lim_{t\to 0}\frac{dt}{ds}=\lim_{t\to 0}\pm\sqrt{\frac{-1}{\left(-1+\frac{1}{(1+\frac{4}{9}|t|^{3})^{2}}\right)|t|}}=\pm\infty.$ (5)
While the components of $\gamma^{\prime}$ do not diverge in the $(T,x)$-coordinate system, both $\frac{dx}{ds}$ and $\frac{dt}{ds}$ diverge in $M_{L}$ in the $(t,x)$-coordinates. Because of this dependency of coordinates the criterion of divergence is not useful for defining pseudo-timelike and pseudo-spacelike curves. That is where the coordinate-independent generalized affine parameter comes into play.
Definition 8 (Pseudo-timelike curve).
Let $M=M_{L}\cup\mathcal{H}\cup M_{R}$ be an $n$-dimensional transverse type-changing singular semi-Riemannian manifold, $g$ a type-changing metric, and $\mathcal{H}:=\{q\in M\!\!:g\!\!\mid_{q}is\;degenerate\}$ the locus of signature change. We assume that one component, $M_{L}$, of $M\setminus\mathcal{H}$ is Lorentzian, and the other component, $M_{R}$, is Riemannian.
Let $\gamma\colon[a,b]\to M$ be a continuous and differentiable curve, with $\subset\mathbb{R}, <a<b<-\infty<a<b$$\infty$. We call $\gamma=\gamma^{\mu}(u)=x^{\mu}(u)$ in $M$ a pseudo-timelike (respectively, pseudo-spacelike) curve if
1. $\operatorname{Im}(\gamma)\cap M_{L}\neq\varnothing$, i.e. $\gamma$ has image points in the Lorentzian region, and
2. for every generalized affine parametrization of $\gamma$ in $M_{L}$ there exists $\varepsilon>0$ such that
$g(\gamma^{\prime},\gamma^{\prime})<-\varepsilon\quad\text{(respectively, }g(\gamma^{\prime},\gamma^{\prime})>\varepsilon\text{)}.$
Example 2.
Revisiting Example 1, we find that both coordinate vector fields, $\frac{\partial}{\partial T}$ and $\frac{\partial}{\partial x}$, are covariantly constant in $M_{L}$ and $M_{R}$ (this is because the Christoffel symbols all vanish in the $(T,x)$-coordinate system). Hence, we can parallel transport $\frac{\partial}{\partial T}$ and $\frac{\partial}{\partial x}$ along any curve in $M_{L}$ and $M_{R}$, with the transport being path-independent (no anholonomy).
Since we aim at parametrizing the curve $\gamma$ by the generalized affine parameter $\mu$ with respect to the coordinate vector fields $\frac{\partial}{\partial T}=\frac{1}{\sqrt{\mid t\mid}}\frac{\partial}{\partial t}$ and $\frac{\partial}{\partial x}$ we are able to start with an arbitrary parametrization. Hence, let $\gamma(t)=(T(t),x(t))$ be parametrized by $t$, and then $\dot{\gamma}(t)=\frac{dT}{dt}\frac{\partial}{\partial T}+\frac{dx}{dt}\frac{\partial}{\partial x}$. By means of Definition 6 we immediately get
$V^{0}(t)=\frac{dT}{dt}=\sqrt{\mid t\mid}$
and
$V^{1}(t)=\frac{dx}{dt}=\frac{d}{dt}\arctan\left(\frac{2}{3}\sqrt{\mid t\mid}^{3}\textrm{sgn}(t)\right).$
In $M_{L}$ this yields
$V^{1}(t)=\frac{\sqrt{\mid t\mid}}{1+\frac{4}{9}\mid t\mid^{3}}.$
Consider now $\tilde{\gamma}(t(s))=\gamma(s)$, in which $\tilde{\gamma}$ is related to the curve $\gamma$ by reparametrization of $\gamma$ by $t$. With this notation we have the basis fields $E_{\tilde{\gamma}(t),0}=\frac{1}{\sqrt{\mid t\mid}}\frac{\partial}{\partial t}$ and $E_{\tilde{\gamma}(t),1}=\frac{\partial}{\partial x}$ along $\tilde{\gamma}$. The reparametrized curve $\tilde{\gamma}(t(s))$ also gives
$\dot{\tilde{\gamma}}(t)=V^{i}(t)E_{\tilde{\gamma}(t),i}=\frac{\partial}{\partial t}+\frac{dx}{dt}\frac{\partial}{\partial x}.$
The Definition 6 for the generalized affine parameter gives
$\frac{d\mu}{dt}=\sqrt{(V^{0}(t))^{2}+(V^{1}(t))^{2}}=\sqrt{\mid t\mid+\frac{\mid t\mid}{(1+\frac{4}{9}\mid t\mid^{3})^{2}}}.$
It now follows easily that for the reparametrization of $\hat{\gamma}(t)$ by the generalized affine parameter $\mu$ (i.e. $\hat{\gamma}(\mu(t))=\tilde{\gamma}(t)$) we have in $M_{L}$:
$g(\dot{\hat{\gamma}}(\mu(t)),\dot{\hat{\gamma}}(\mu(t)))=g(\frac{d\hat{\gamma}(\mu(t))}{d\mu},\frac{d\hat{\gamma}(\mu(t))}{d\mu})=g(\frac{1}{\frac{d\mu}{dt}}\dot{\hat{\gamma}}(t),\frac{1}{\frac{d\mu}{dt}}\dot{\hat{\gamma}}(t))$
$\frac{g(\frac{\partial}{\partial t}+\frac{dx}{dt}\frac{\partial}{\partial x},\frac{\partial}{\partial t}+\frac{dx}{dt}\frac{\partial}{\partial x})}{(\frac{d\mu}{dt})^{2}}=\frac{t+(\frac{dx}{dt})^{2}}{\mid t\mid+\frac{\mid t\mid}{(1+\frac{4}{9}\mid t\mid^{3})^{2}}}=\frac{t+\frac{\mid t\mid}{(1+\frac{4}{9}\mid t\mid^{3})^{2}}}{\mid t\mid+\frac{\mid t\mid}{(1+\frac{4}{9}\mid t\mid^{3})^{2}}}.$
Taking the limit
$\underset{t\rightarrow 0^{-}}{\lim}\frac{t+\frac{\mid t\mid}{(1+\frac{4}{9}\mid t\mid^{3})^{2}}}{\mid t\mid+\frac{\mid t\mid}{(1+\frac{4}{9}\mid t\mid^{3})^{2}}}=0$
reveals that the curve $\gamma$ is not pseudo-timelike as it does not meet the $\varepsilon$-requirement of Definition 8.
In Section 2.1 we repeatedly rather vaguely referred to the concept of a timelike (or spacelike, respectively) curve that asymptotically becomes lightlike. The above example highlights how the notion of “asymptotically lightlike” should be understood. A timelike (or spacelike, respectively) curve in $M_{L}$ that is not pseudo-timelike (or pseudo-spacelike, respectively) can be thus specified as asymptotically lightlike.
Example 3.
Finally, if we modify the previously discussed curve $\gamma$ by keeping the $t$-coordinate but stating $x=0$, we get a curve $\alpha$. With the same notation as above, we then get $V^{0}(t)=\sqrt{\mid t\mid}$, $V^{1}(t)=0$ and $\frac{d\mu}{dt}=\sqrt{\mid t\mid}$. Hence, this results in $g(\dot{\hat{\alpha}}(\mu(t)),\dot{\hat{\alpha}}(\mu(t)))=\frac{1}{\mid t\mid}g(\frac{\partial}{\partial t},\frac{\partial}{\partial t})=\frac{t}{\mid t\mid}=-1$ in the Lorentzian region $M_{L}$. The curve $\alpha$ is pseudo-timelike as it obviously does meet the $\varepsilon$-requirement of Definition 8.
This disquisition makes it clear why the notion of the generalized affine parameter is necessary and useful in order to define pseudo-timelike and pseudo-spacelike curves. If we were to loosen the requirement in Definition 8 by replacing “for every generalized affine parametrization of $\gamma$ in $M_{L}$” with “for every affine parametrization of $\gamma$ in $M_{L}$”, then no curve that is timelike in the Lorentz sector and reaches the hypersurface $\mathcal{H}$ would be pseudo-timelike throughout the entire manifold $M$. (However, this statement applies only to curves that actually reach the hypersurface $\mathcal{H}$. Timelike curves that lie entirely within $M_{L}$ and maintain a “distance” from $\mathcal{H}$ due to a tubular neighborhood within $M_{L}$ also satisfy the relaxed condition, as they only have affine parameters with $g(\gamma^{\prime},\gamma^{\prime})=\text{const}<0$.)
Similarly, any timelike curve in $M_{L}$ would meet the requirements of a pseudo-timelike curve if we modified the definition by requesting “for a suitable parametrization of $\gamma$ in $M_{L}$” instead of “for every generalized affine parametrization of $\gamma$ in $M_{L}$”. In this regard, the concept of the generalized affine parameter is the right tool to tell apart suitable from unsuitable curves for the definition of pseudo-timelike and pseudo-spacelike curves.
Interestingly, our rationale for the new definition of a pseudo-timelike curve is reminiscent of the analysis undertaken in [30]. In Section 2 of [30] the distinction between causal curves, timelike almost everywhere curves and timelike curves is introduced in which the latter one is defined as follows: A timelike curve is a causal curve $\gamma\colon I\longrightarrow M$ such that $g(\gamma^{\prime},\gamma^{\prime})<-\varepsilon$ almost everywhere for some $\varepsilon>0$.
The author illustrates the situation in his Figure 1 which contrasts a timelike curve with a timelike almost everywhere curve. The latter one can not be viewed as a timelike curve because it approaches a null vector at its break point. Compared to our setting, however, the culprit here is that the curve is not differentiable at the breaking point. However, if we were to make the curve differentiable by bending its upper section, it still wouldn’t be timelike. Its restriction to the lower region before the inflection point is timelike, but it cannot be extended upwards into a timelike curve. Now, imagine we are not in Minkowski spacetime, but instead, a signature-type change occurs at the (former) inflection point so that the “upper half” of the space becomes Euclidean (in this case, the figure would correspond to the $(T,x)$-coordinates, not the $(t,x)$-coordinates, in the toy model). In this scenario, the curve restricted to the Lorentz sector would be timelike, but after the signature change is “reversed”, it cannot be extended upwards into a timelike curve. In this sense, the entire curve in the signature-changing version of this example is not pseudo-timelike.
Remark 3.
The introduction of pseudo-timelike curves is not merely a technical tool. If a timelike trajectory were to become asymptotically null near the degenerate hypersurface, this would correspond to a massive particle being accelerated to arbitrarily close to the speed of light. Such behaviour would require a divergent energy input and is therefore unphysical. Pseudo-timelike curves exclude this scenario and ensure that the worldlines of massive particles remain genuinely timelike.
Now we can slightly modify the definition of a (simply) closed curve in order for it to correctly apply to signature-type changing singular semi-Riemannian manifolds $M$ with a metric $g$:
Definition 9 (Chronology-violating curve).
A smooth, pseudo-timelike curve $\gamma\colon I\longrightarrow M$ is said to be chronology-violating if there exists a subset of $\gamma[I]$ homeomorphic to $S^{1}$ such that there are at least two parameters $s_{1},s_{2}\in I$ with $\gamma(s_{1})=\gamma(s_{2})$. Moreover, $\gamma$ must belong to one of the following two classes: [^7] Note that this means that there must be at least one such subset to fulfill this definition.
  1. 1.
    The pseudo-timelike curve $\gamma$ is periodic, i.e the image $\gamma[I]$ is homeomorphic to $S^{1}$. Moreover, for $s_{1}$, $s_{2}\in I$ the associated tangent vectors, $\gamma^{\prime}(s_{1})$ and $\gamma^{\prime}(s_{2})$, are timelike and positively proportional. We denote this type of curve as closed pseudo-timelike curve.
  2. 2.
    The curve $\gamma$ intersects itself for $s_{1}$, $s_{2}\in I$ and the associated tangent vectors, $\gamma^{\prime}(s_{1})$ and $\gamma^{\prime}(s_{2})$, are timelike whereas the tangent directions are not necessarily the same (i.e. they do not need to be positively proportional). This type of curve is said to contain a loop.
3 Global structure of signature-type changing semi-Riemannian manifolds
First, let us revisit the definitions of the following concepts related to manifold orientability.
Definition 10.
[7] A smooth $n$-dimensional manifold $M$ is orientable if and only if it has a smooth global nowhere vanishing $n$-form (also called a top-ranked form). [^8] An orientation of $M$ is the choice of a continuous pointwise orientation, i.e. the specific choice of a global nowhere vanishing $n$-form.
For a differentiable manifold to be orientable all that counts is that it admits a global top-ranked form - it is not important which specific top-ranked form is selected.
To ensure thoroughness, we also want to mention the definition of parallelizability, which likewise does not involve any metric and is therefore again applicable to manifolds with changing signature types. It is well-known that a manifold $M$ of dimension $n$ is defined to be parallelizable if there are $n$ global vector fields that are linearly independent at each point. We define it similarly to the approach in [8]:
Definition 11.
A smooth $n$-dimensional manifold $M$ is parallelizable if there exists a set of smooth vector fields $\{V,E_{1},\ldots,E_{n-1}\}$ on $M$, such that at every point $p\in M$ the tangent vectors $\{V(p),E_{1}(p),\ldots,E_{n-1}(p)\}$ provide a basis of the tangent space $T_{p}M$. A specific choice of such a basis of vector fields on $M$ is called an absolute parallelism of $M$.
Equivalently, a manifold $M$ of dimension $n$ is parallelizable if its tangent bundle $TM$ is a trivial bundle, so that the associated principal bundle of linear frames has a global section on $M$, i.e. the tangent bundle is then globally of the form $TM\simeq M\times\mathbb{R}^{n}$. Moreover, it is worth pointing out that orientability and also parallelizability are differential topological properties which do not depend on the metric structure, but only on the topological manifold with a globally defined differential structure.
Remark 4.
It is worth mentioning that given an absolute parallelism of $M$, one can use these $n$ vector fields to define a basis of the tangent space at each point of $M$ and thus one can always get a frame-dependent metric $g$ by defining the frame to be orthonormal. Moreover, the special orthogonal group, denoted $SO(n,\mathbb{R})$, acts naturally on each tangent space via a change of basis, it is then possible to obtain the set of all orthonormal frames for $M$ at each point qua the oriented orthonormal frame bundle of $M$, denoted $F_{SO}(M)$, associated to the tangent bundle of $M$.
The next three definitions, however, depend not only on the underlying manifold but also on its specific type-changing metric $g$. For our purpose, let $(M,g)$ be a smooth, signature-type changing manifold (possibly with boundary).
Definition 12 (Pseudo-timelike).
A vector field $V$ on a signature-type changing manifold $(M,g)$ is pseudo-timelike if and only if its integral curves are pseudo-timelike (hence, in particular, $V$ is timelike in $M_{L}$). [^9] Keep in mind that a timelike vector field is a vector field $V$ on a spacetime manifold $(M,g)$ where the vectors at every point are timelike, meaning $g(V(p),V(p))<0$ for all points $p$ on the manifold.
Definition 13 (Pseudo-time orientable).
A signature-type changing manifold $(M,g)$ is pseudo-time orientable if and only if the Lorentzian region $M_{L}$ is time orientable. [^10] A pseudo-time orientation of such a manifold $(M,g)$ corresponds to the specific choice of a continuous non-vanishing pseudo-timelike vector field $V$ on $M$.
Lemma 1.
A singular semi-Riemannian manifold $(M,g)$ is pseudo-time orientable if and only if there exists a vector field $X\in\mathfrak{X}(M)$ that is pseudo-timelike.
Proof.
$\Longrightarrow"$ Let a singular semi-Riemannian manifold $(M,g)$ be pseudo-time orientable. This means the Lorentzian region $M_{L}$ is time orientable. A Lorentzian manifold is time-orientable if there exists a continuous timelike vector field. Accordingly, there must exist a continuous timelike vector field $X\in\mathfrak{X}(M_{L})$ in the Lorentzian region. As per Definition 12, a vector field $X$ in a signature-type changing manifold is pseudo-timelike if and only if $X$ is timelike in $M_{L}$ and its integral curve is pseudo-timelike; this means that $X$ is allowed to vanish on $M_{R}$. [^11] In this part of the proof, the only thing that matters is whether the “pseudo-timelike vector field” is allowed to vanish on $M_{R}$. This question is independent of whether the “generalized affine parameter” condition is required in $M_{L}$, because the issue of whether the vector field “is allowed to vanish on $M_{R}$” concerns only its “magnitude”, while the “generalized affine parameter” condition pertains solely to its “direction” (specifically, that the vector field is not asymptotically lightlike). Hence, we can extend the vector field $X$ arbitrarily to all of $M$, and per definition $X\in\mathfrak{X}(M)$ is pseudo-timelike.
[…]
According to that, such a definition of a pseudo-time orientation is possible if $M_{L}$ admits a globally consistent sense of time, i.e. if in $M_{L}$ we can continuously define a division of non-spacelike vectors into two classes. For a transverse, signature-type changing manifold (with a transverse radical), this definition arises naturally because, in $M_{R}$, all vectors can be considered spacelike. Additionally, all non-spacelike vectors on $\mathcal{H}$ are lightlike. [^12] Note that this applies generally, including in the case of a tangent radical, since there are no timelike vectors on $\mathcal{H}$. However, the subsequent division into two classes requires a transverse radical. In the case that $Rad_{q}\cap T_{q}\mathcal{H}=\{0\}$ $\forall q\in\mathcal{H}$, these lightlike vectors can be naturally divided into two classes: those pointing towards $M_{L}$ and those pointing towards $M_{R}$.
Example 4.
Figure 3: Riemannian and Lorentzian region in the Hartle-Hawking no-boundary model.
Consider the classic type of a spacetime $M$ with signature-type change which is obtained by cutting an $S^{4}$ along its equator and joining it to the corresponding half of a de Sitter space, Figure 3. The de Sitter spacetime is time-orientable [32], hence $M$ is pseudo-time orientable.
Definition 14 (Pseudo-space orientable).
A signature-type changing manifold $(M,g)$ of dimension $n$ is pseudo-space orientable if and only if it admits a continuous non-vanishing spacelike (n1)(n-1)-frame field on $M_{L}$. This is a set of n1n-1 pointwise orthonormal spacelike vector fields on $M_{L}$. [^13] A pseudo-space orientation of a manifold $(M,g)$ corresponds to the specific choice of a continuous non-vanishing field of orthonormal spacelike (n1)(n-1)-beins on $M_{L}$.
Proposition 5.
[29] Every parallelizable manifold $M$ is orientable.
In Lorentzian geometry the fact of $M$ being time-orientable and space-orientable implies that $M$ is orientable [20]. The proposition below illustrates that this result from Lorentzian geometry cannot be applied to signature-type changing manifolds.
Proposition 6.
Even if a transverse, signature-type changing manifold $(M,g)$ with a transverse radical is pseudo-time orientable and pseudo-space orientable, it is not necessarily orientable.
Proof.
Consider an arbitrary manifold of $\dim(M)=2$ with a change of signature, for which the conditions of Proposition 6 are given (in higher dimensions, the same idea can be carried out through a trivial augmentation of dimensions). In case this manifold is non-orientable, there is nothing to show. However, if it is orientable, cut out a disk from the Riemannian sector and replace it with a crosscap, equipped with any Riemannian metric. In a tubular neighborhood of the cutting line, construct a Riemannian metric that mediates between the metrics of the crosscap and the rest (this is possible due to the convexity of the space formed by all Riemannian metrics). This surgical intervention results in the transition to a non-orientable manifold with a change of signature. Since the intervention is limited to the Riemannian sector, the conditions of the proposition remain unaffected. Thus Proposition 6 is proven. ∎
Remark 5.
One can always “switch” between non-orientability and orientability using the crosscap. Starting with an orientable manifold, one transitions to non-orientable by replacing a crosscap (if already present) with a disk. If no crosscap is present, such a transition occurs by replacing a disk with a crosscap.
Example 5.
The Möbius strip $\mathbb{M}$ has a non-trivial vector bundle structure over $S^{1}$, which means that the bundle cannot be trivialized globally. Specifically, the Möbius strip is a line bundle over $S^{1}$ with a non-trivial twist. [^14] The Möbius strip is particularly interesting because it can be found on any arbitrary non-orientable surface. Additionally, any Lorentzian manifold $\mathbb{M}\times\mathbb{R}^{n}$ based on the Möbius strip crossed with $\mathbb{R}^{n}$ either fails to be time-orientable or space-orientable [16]. Hence, $\mathbb{M}$ is neither parallelizable nor orientable.
To see this, consider the Möbius strip $\mathbb{M}=\mathbb{R}\times\mathbb{R}/\sim$ with the identification $(t,x)\sim(\tilde{t},\tilde{x})\Longleftrightarrow(\tilde{t},\tilde{x})=((-1)^{k}t,x+k),k\in\mathbb{Z}$. Notice that the identification has no bearing on proper subsets of $((-1)^{k}t,x+k),k\in\mathbb{Z}$, and the fibre $\mathbb{R}$ is a vector space.
As $\mathbb{M}$ is a fiber bundle over the base space $S^{1}$, a section of that fiber bundle must be a continuous map $\sigma:S^{1}\longrightarrow\mathbb{M}$ such that $\sigma(x)=(h(x),x)\in\mathbb{M}$. For $\sigma$ to be continuous, $h$ must satisfy h(0)=h(k)-h(0)=h(k). The intermediate value theorem guarantees that there is some $\tilde{x}\in[0,k]$ such that $h(\tilde{x})=0$. This means that every section of $\mathbb{M}$ intersects the zero section, and the sections that form a basis for the fibre are not non-zero everywhere.
Definition 15.
A pseudo-spacetime is a $4$-dimensional, pseudo-time oriented, semi-Riemannian manifold with a type-changing metric.
Proposition 7.
Let $(\mathbb{R}^{n},g)$ be a transverse, signature-type changing $n$-manifold with a transverse radical, and let $\mathcal{H}\subset\mathbb{R}^{n}$ be a codimension one closed hypersurface of signature change without boundary. [^15] Here “closed” is meant in the topological sense of “the complement of an open subset of $\mathbb{R}^{n}$” and not in the manifold sense of “a manifold without boundary that is compact.” Then $\mathcal{H}$ is always orientable.
Proof.
This can be shown by a purely topological argument, as in [33]. ∎
Proposition 8.
Let $(M,g)$ be a transverse, signature-type changing, oriented, $n$-dimensional manifold with a transverse radical, and let $\mathcal{H}\subset M$ be the hypersurface of signature change. Then $\mathcal{H}$ is also oriented.
Proof.
The hypersurface of signature change, as a closed submanifold of codimension one, is the inverse image of a regular value of a smooth map $f\colon M\to\mathbb{R}$. Specifically, $\mathcal{H}=f^{-1}(c)$ for some regular value $c\in\mathbb{R}$. The manifold $M$ is oriented, so its tangent bundle $TM$ is oriented, meaning there is a consistent choice of orientation on each tangent space $T_{p}M$ for $p\in M$. Since $\mathcal{H}$ is a hypersurface in $M$, at each point $q\in\mathcal{H}$, the tangent space $T_{q}\mathcal{H}$ is a subspace of the tangent space $T_{q}M$ of dimension n1n-1, and therefore $T\mathcal{H}$ is a subbundle of $TM$. The remaining direction in $T_{q}M$ can be described by a normal vector $N(q)$, which is a vector in $T_{q}M$ that is perpendicular to $T_{q}\mathcal{H}$.
Since $M$ is oriented, for each point $q\in\mathcal{H}$, the tangent space $T_{q}M$ has an orientation that can be described by an ordered basis, say $\{v_{1},\ldots,v_{n-1},N(q)\}$, where $\{v_{1},\ldots,v_{n-1}\}$ is an oriented basis for $T_{q}\mathcal{H}$ and $N(q)$ is the normal vector. Hence, this induces a consistent orientation on $T_{q}\mathcal{H}$ across all points $q\in\mathcal{H}$, since the orientation of $M$ provides a consistent choice of $N(q)$ across $\mathcal{H}$. Therefore, $\mathcal{H}$ inherits a consistent orientation from $M$, proving that $\mathcal{H}$ is oriented.
Moreover, without loss of generality, we can choose $1$ as a regular value (see also [23]). Thus, $\mathcal{H}:=f^{-1}(1)=\{p\in M\mid f(p)=1\}$ is a submanifold of $M$ of dimension n1n-1. For every $q\in\mathcal{H}$, the tangent space $T_{q}\mathcal{H}=T_{q}(f^{-1}(1))$ to $\mathcal{H}$ at $q$ is the kernel $\ker(df_{q})$ of the map $df_{q}\colon T_{q}M\to T_{1}\mathbb{R}$. Then $T_{q}\mathcal{H}=\left\langle\text{grad}f_{q}\right\rangle^{\bot}$, and therefore the gradient $\text{grad}f$ yields an orientation of $\mathcal{H}$. ∎
Provided a transverse, signature-type changing manifold $(M,g)$ with a transverse radical is pseudo-time orientable, then we can choose one of the two possible time orientations at any point in each connected component of $M_{L}$, and thus designating the future direction of time in the Lorentzian regime. On $\mathcal{H}$ all non-spacelike vectors are lightlike and smoothly divided into two classes in a natural way: the vectors located at an initial base point on $\mathcal{H}$ are either pointing towards $M_{L}$ or towards $M_{R}$. This together with the existent absolute time function (that establishes a time concept [25] in the Riemannian region) can be considered as arrow of time on $M$.
Definition 16 (Natural time direction).
Let $(M,g)$ be a pseudo-time orientable, transverse, signature-type changing $n$-dimensional manifold with a transverse radical. Then in the neighborhood of $\mathcal{H}$ the absolute time function $\mathfrak{h}(t,\mathbf{\hat{x}}):=t$, where $\mathbf{\hat{x}}=(x^{1},\ldots,x^{n-1})$, imposes a natural time direction by postulating that the future corresponds to the increase of the absolute time function. In this way, the time orientation is determined in $M_{L}$.
Remark 6.
[…]
Figure 4: In the left example the curves $\alpha$ and $\gamma$ are both future-directed. The curve $\beta$ runs within the edge that is twisted and identified with the left edge; therefore $\beta$ is neither future-directed nor past-directed. In the right example the curves $\alpha$, $\beta$ and $\gamma$ are future-directed. In both examples the loops around $\mathcal{H}$ are neither future-directed nor past-directed.
Definition 17 (Future-directed).
A pseudo-timelike curve (see Definition 8) in $(M,g)$ is future-directed (in the sense of Definition 16 and Remark 6) if for every point in the curve:
  1. (i)
    within $M_{L}$ the tangent vector is future-pointing, and
  2. (ii)
    on $\mathcal{H}$ the associated tangent vector with an initial base point on $\mathcal{H}$ is future-pointing.
Respective past-directed curves are defined analogously. Notice that, per assumption, one connected component of $M\setminus\mathcal{H}$ is Riemannian and all other connected components $(M_{L_{\alpha}})_{\alpha\in I}\subseteq M_{L}\subset M$ are Lorentzian. This configuration could (at least locally) potentially allow for a $M_{L}-M_{R}-M_{L}$-sandwich-like structure of $M$, where $\mathcal{H}$ consists of two connected components $(\mathcal{H}_{\alpha})_{\alpha\in\{1,2\}}$. Consequently, this would also imply the existence of two absolute time functions, see Figure 4.
Definition 18 (Pseudo-chronological past and future).
Let $(M,g)$ be a pseudo-time orientable, transverse, signature-type changing $n$-dimensional manifold with a transverse radical.
$\mathcal{I}^{-}(p)=\{q\in M\colon q\ll p\}$ is the pseudo-chronological past of the event $p\in M$. In other words, for any two points $q,p\in M$, we write $q\ll p$ if there is a future-directed pseudo-timelike curve from $q$ to $p$ in $M$.
$\mathcal{I}^{+}(p)=\{q\in M\colon p\ll q\}$ is the pseudo-chronological future of the event $p\in M$. In other words, for any two points $p,q\in M$, we write $p\ll q$ if there is a future-directed pseudo-timelike curve from $p$ to $q$ in $M$.
Figure 5: For an event $p\in\mathcal{H}$ there exists a future-directed pseudo-timelike curve (as depicted) that connects the points $p$ and $q$ in $M$. Similarly any point in $M$ can be reached by such a future-directed pseudo-timelike curve from $p$. That is why for the pseudo-chronological future we have $\mathcal{I}^{+}(p)=\{q\in M\colon p\ll q\}=M$.
Remark 7.
Interestingly, this definition leads to the following peculiar situation: Recall that any curve is denoted pseudo-timelike if its $M_{L}$-segment is timelike. To that effect, all curves that steer clear of $M_{L}$ (and do not have a $M_{L}$–segment) are also considered pseudo-timelike. When $p\in\mathcal{H}\cup M_{R}$ then the pseudo-chronological past of $p$ is $\mathcal{I}^{-}(p)=M\setminus M_{L}$ and the pseudo-chronological future of $p$ is $\mathcal{I}^{+}(p)=M$, see Figure 5.
4 Chronology violating pseudo-timelike loops
In Section 2, we introduced the notion of closed pseudo-timelike curves on a signature-type changing background and we demonstrated how they must be defined to ensure that the concept of causality remains meaningful. In this section, we will reveal the non-well-behaved nature of transverse, signature-type changing, $n$-dimensional manifolds with a transverse radical.
4.1 Local pseudo-timelike loops
In a sufficiently small region near the junction of signature change, these manifolds exhibit local anomalies. Specifically, each point on the junction gives rise to the existence of closed time-reversing loops, challenging conventional notions of temporal consistency. One of our main results, Theorem 2, can now be proved quite easily.
Theorem (Local loops).
[…]
Proof.
Let $\tilde{g}=-t(dt)^{2}+\tilde{g}_{jk}(t,x^{1},\ldots,x^{n-1})dx^{i}dx^{k}$, $j,k\in\{1,\ldots,n-1\}$, be a transverse, signature-type changing metric with respect to a radical-adapted Gauss-like coordinate patch $(U_{\varphi},\varphi)$ with $U_{\varphi}\cap\mathcal{H}\neq\emptyset$. [^16] This is, $U_{\varphi}$ is sufficiently small to be expressed in the adapted radical-adapted Gauss-like coordinate system $\xi(U_{\varphi})$. Choose smooth coordinates $(t_{0},x_{0}^{1},\ldots,x_{0}^{n-1})$ with $t_{0}>0$ and $\xi_{0}>0$, such that
$C_{0}:=[0,t_{0}]\times B_{\xi_{0}}^{n-1}=[0,t_{0}]\times\{x\in\mathbb{R}^{n-1}\mid{\displaystyle\sum_{k=1}^{n-1}(x^{k})^{2}\leq\xi_{0}^{2}\}\subset}\mathbb{R}^{n}$
is contained in the domain of the coordinate chart (open neighborhood) $U_{\varphi}$. Then
$C_{0}\times\mathbb{S}^{n-2}=C_{0}\times\{v\in\mathbb{R}^{n-1}\mid{\displaystyle\sum_{k=1}^{n-1}(v^{k})^{2}=1\}}$
as a product of two compact sets is again compact.
Next, consider the function
$\tilde{G}\colon C_{0}\times\mathbb{S}^{n-2}\longrightarrow\mathbb{R},$
$(t,x^{1},\ldots,x^{n-1},v^{1},\ldots,v^{n-1})\mapsto\tilde{g}_{jk}(t,x^{1},\ldots,x^{n-1})v^{j}v^{k}.$
As $\tilde{G}$ is a smooth function defined on the compact domain $C_{0}\times\mathbb{S}^{n-2}$, by the Extreme Value Theorem it has an absolute minimum $G_{0}$. Hence, on $(U_{\varphi},\varphi)$ we can uniquely define $\tilde{g}_{0}=-t(dt)^{2}+G_{0}\delta_{jk}dx^{j}dx^{k},\>j,k\in\{1,\ldots,n-1\}$.
By this definition, for all nonzero lightlike vectors $X\in T_{p}M,\>p\in C_{0}$ with respect to $\tilde{g}_{0}$, we have $\tilde{g}_{0}=-t(X^{0})^{2}+G_{0}\delta_{jk}X^{j}X^{k}=0\Longleftrightarrow-t(X^{0})^{2}=-G_{0}\delta_{jk}X^{j}X^{k}$, then
$\tilde{g}(X,X)=-t(X^{0})^{2}+\tilde{g}_{jk}(t,x^{1},\ldots,x^{n-1})X^{j}X^{k}$
$G_{0}\delta_{jk}X^{j}X^{k}+\tilde{g}_{jk}(t,x^{1},\ldots,x^{n-1})X^{j}X^{k}$
$\delta_{jk}X^{j}X^{k}\cdot(-G_{0}+\tilde{g}_{rs}(t,x^{1},\ldots,x^{n-1})\frac{X^{r}}{\sqrt{\delta_{ab}X^{a}X^{b}}}\frac{X^{s}}{\sqrt{\delta_{cd}X^{c}X^{d}}})\geq 0.$
Clearly, $\tilde{g}(X,X)\geq 0$ because $G_{0}>0$ per definition and $\delta_{jk}X^{j}X^{k}=\frac{t(X^{0})^{2}}{G_{0}}\geq 0$. Therefore, the vector $X\in T_{p}M,\>p\in C_{0}$ is not timelike with respect to $\tilde{g}$. This means, within $C_{0}$ the $\tilde{g}-$light cones always reside inside of the $\tilde{g}_{0}$-light cones, i.e. $\tilde{g}\leq\tilde{g}_{0}$ in $C_{0}$. The cull cones of $\tilde{g}_{0}$ are more opened out than those of the metric $\tilde{g}$. Denote $p_{0}\in C_{0}$ by $(t(p_{0}),x^{1}(p_{0}),\ldots,x^{n-1}(p_{0}))=(t_{0},x_{0}^{1},\ldots,x_{0}^{n-1})$.
[…]
Figure 6: The chronological past $I^{-}(p_{1})$ of a point $p_{1}\in U_{\varphi}$.
For a lightlike curve $\alpha(t)\colon I\longrightarrow U_{\varphi}$ with starting point $p_{0}$, we have $\delta_{jk}\frac{dx^{j}}{dt}\frac{dx^{k}}{dt}>0$ for each slice $(U_{\varphi})_{t_{i}}$ with $t\neq 0$. Lightlike curves with starting point $p_{0}$ can be parametrized with the Euclidean arc length $\sigma$ in $B_{\xi_{0}}^{n-1}$, such that $(\tilde{g}_{0})_{t}(\dot{\alpha}(\sigma),\dot{\alpha}(\sigma))=\delta_{jk}\frac{dx^{j}}{d\sigma}\frac{dx^{k}}{d\sigma}=1$, $\forall\>\sigma\in I$, where $I$ is some interval in $\mathbb{R}$. More precisely, $\sigma$ can be considered as arc length (parameter) in terms of some auxiliary Riemannian metrics, each defined on a hypersurface with $t=\textrm{const}$. Consequently we get
$0=\tilde{g}_{0}(\dot{\alpha}(\sigma),\dot{\alpha}(\sigma))=-t(\dot{\alpha}^{0})^{2}+G_{0}\delta_{ik}\dot{\alpha}^{j}\dot{\alpha}^{k}$
$t(\frac{dt}{d\sigma})^{2}+G_{0}\underset{1}{\underbrace{\delta_{ik}\frac{dx^{j}}{d\sigma}\frac{dx^{k}}{d\sigma}}=}-t(\frac{dt}{d\sigma})^{2}+G_{0},$
and this implies
$\frac{d\sigma}{dt}=\pm\sqrt{\frac{t}{G_{0}}}\Longrightarrow\sigma(t)=\pm\int\sqrt{\frac{t}{G_{0}}}dt=\pm\frac{2}{3}t\sqrt{\frac{t}{G_{0}}}+\textrm{const}.$
Since $\sigma$ is given as a function of $t$, it represents the arc length from the starting point at $t(p_{0})=t_{0}$ to $t(0)=0$. Then past-directed $\tilde{g}_{0}$-lightlike curves emanating from $p_{0}$ reach the hypersurface at $t=0$ after passing through the arc length distance
$\triangle\sigma=\pm{\displaystyle\int_{0}^{t_{0}}}\sqrt{\frac{t}{G_{0}}}dt=\pm\frac{2}{3}t_{0}\sqrt{\frac{t_{0}}{G_{0}}}+\textrm{const}.=\pm\frac{2}{3}\sqrt{\frac{t_{0}^{3}}{G_{0}}}+\textrm{const}$
along the said section of the curve from the fixed starting point $p_{0}$.
Provided this arc length distance satisfies $\triangle\sigma\leq\xi_{0}$, then the past-directed lightlike curves $\alpha(t)$ (emanating from $p_{0}$) reach the hypersurface at $t=0$ while remaining within $C_{0}$. Accordingly this is also the case for $\tilde{g}-$lightlike curves emanating from $p_{0}$. Conversely, if $\triangle\sigma>\xi_{0}$ then there exist past-directed $\tilde{g}_{0}$-lightlike curves emanating from $p_{0}$ that reach the hypersurface outside of $C_{0}$.
In this case, we have $\triangle\sigma=\frac{2}{3}\sqrt{\frac{t_{0}^{3}}{G_{0}}}>\xi_{0}\Longleftrightarrow t_{0}>\sqrt[3]{\frac{9}{4}\xi_{0}^{2}\cdot G_{0}}$ and we must adjust the new starting point $p_{1}=(t_{1},x_{0})$ accordingly by setting $t_{1}\leq\sqrt[3]{\frac{9}{4}\xi_{0}^{2}\cdot G_{0}}<t_{0}$. Thereby we make sure that all past-directed $\tilde{g}-$lightlike curves emanating from $p_{1}$ hit the hypersurface $\mathcal{H}$ without leaving $C_{0}$. That is $\overline{I_{\tilde{g}}^{-}(p_{1})}\subset C_{0}\subset U_{\varphi}\subset M$, where $\overline{I_{\tilde{g}}^{-}(p_{1})}$ is the $\tilde{g}-$chronological past of the event $p_{1}\in M_{L}$, restricted to $M_{L}\cup\mathcal{H}$, see Figure 6.
It now suffices to connect two of such points $\hat{x}_{1},\hat{x}_{2}\in\overline{I_{\tilde{g}}^{-}(p_{0})}\cap\mathcal{H}$ (or, if need be $\overline{I_{\tilde{g}}^{-}(p_{1})}\cap\mathcal{H}$) in an arbitrary fashion within the Riemannian sector $M_{R}$. By what a pseudo-timelike loop gets generated, if $U_{\varphi}$ was chosen small enough. ∎
Summarized, for each neighborhood $U(q)$ that admits radical-adapted Gauss-like coordinates $\xi=(t,\mathbf{\hat{x}})=(t,x^{1},\ldots,x^{n-1})$ centered at some $q\in\mathcal{H}$, and $U(q)\cap\mathcal{H}\neq\emptyset$, we are able to pick a point $p_{0}\in U(q)$ and an associated compact set $C_{0}\subset U(q)$. For the metric $\tilde{g}$ there exists a corresponding uniquely (i.e., only dependent on the chosen set $C_{0}$) defined metric $\tilde{g}_{0}$ with $\tilde{g}\leq\tilde{g}_{0}$ within $C_{0}$. [^18] The set $C_{0}$ does not need to be “maximal” (in some sense) and is therefore not unique. Then we must distinguish between two cases, that is
i)  with respect to the metric $\tilde{g}_{0}$ we have $I_{0}^{-}(p_{0})\subset C_{0}$, then also $I^{-}(p_{0})\subset C_{0}$ with respect to $\tilde{g}$,
ii) with respect to the metric $\tilde{g}_{0}$ we have the situation $I_{0}^{-}(p_{0})\nsubseteq C_{0}$, then there exists a point $p_{1}=(t_{1},x_{0})\in C_{0}\setminus\mathcal{H}$ with $t_{1}<t_{0}$, such that $I_{0}^{-}(p_{1})\subset C_{0}$, hence also $I^{-}(p_{1})\subset C_{0}$ with respect to $\tilde{g}$.
Thus, for any point $q\in\mathcal{H}$ we can find a sufficiently small neighborhood $\tilde{U}\subset U(q)$ containing a point $p\in M_{L}$, such that all past-directed, causal curves emanating from that point, reach the hypersurface within a sufficiently small set $C_{0}$.
Corollary 10.
[…]
The above corollary follows directly from Theorem 2 because we have proven that all past-directed $\tilde{g}-$lightlike curves emanating from $p_{1}$ hit the hypersurface $\mathcal{H}$ without leaving $C_{0}$. That is $\overline{I_{\tilde{g}}^{-}(p_{1})}\subset C_{0}\subset U_{\varphi}\subset M$, where $\overline{I_{\tilde{g}}^{-}(p_{1})}$ is the closure of the $\tilde{g}-$chronological past of the event $p_{1}\in M_{L}$. Hence, we have also shown that the causal past is within $C_{0}$, and furthermore, $J_{\tilde{g}}^{-}(p_{0})\subset\overline{I_{\tilde{g}}^{-}(p_{0})}$.
And since in every neighborhood of each point $q\in\mathcal{H}$ there always exists a pseudo-timelike loop, we can straightforwardly assert the following
Corollary 11.
[…]
Remark 8.
We note that the loops constructed in the proof of Theorem 2 are entirely contained within the small neighborhoods $U(q)$ and do not extend globally. More generally, pseudo-timelike loops may exist that traverse larger portions of the manifold, particularly if the hypersurface $\mathcal{H}$ has a toroidal or otherwise nontrivial global structure. Allowing self-intersections does not conflict with treating the objects as parameterized curves.
As a matter of course, in the Lorentizan region the tangent space at each point is isometric to Minkowski space which is time orientable. Hence, a Lorentzian manifold is always infinitesimally time- and space-orientable, and a continuous designation of future-directed and past-directed for non-spacelike vectors can be made (infinitesimally and therefore, by continuity, also locally). [^19] In case the Lorentzian manifold is time-orientable, a continuous designation of future-directed and past-directed for non-spacelike vectors can be made allover.
Having said that, the infinitesimal properties of a manifold with a signature change are identical to those of a Lorentzian manifold only within the Lorentzian sector. However, when examining the Riemannian sector and the hypersurface, specific distinctions arise. The Riemannian sector and the hypersurface are not infinitesimally modelable by a Minkowski space. While the Riemannian sector reveals an absence of a meaningful differentiation between past- and future-directed vectors, on the hypersurface, one has the flexibility to make arbitrary assignments of such distinctions at the infinitesimal level. If one now determines on the hypersurface whether the direction towards the Lorentzian sector is the future or past direction, it is not only a reference to the tangent space at a point. Rather, it is a local consideration.
[…]
An observer in the region $M_{L}$ near $\mathcal{H}$ perceives these locally closed time-reversing loops (Figure 7) as the creation of a particle and an antiparticle at two different points $\hat{q},q\in\mathcal{H}$. [^20] Such locally closed time-reversing loops around $\mathcal{H}$ obviously do not satisfy the causal relation $\ll$ as introduced above. This could be taken as an object entering the Riemannian region, then resurfacing in the Lorentzian region and proceeding to propagate “forward in time” with respect to the time orientation in $M_{L}$. Since the mathematical definition of the loop itself does not privilege a direction of traversal, this picture is symmetric.
[…]
Figure 7: A closed time-reversing loop.
For fields, take the conjugate $\psi_{t}^{A}=e^{-i\hat{H}t}\psi^{*}$ of $\psi_{t}=e^{i\hat{H}t}\psi$: The unitary temporal evolution of the field operator for antiparticles arises from the temporal evolution of the field operator for particles by applying the same Hamiltonian operator to the adjoint field operator under time reversal. Some literature [18] points to the idea that concepts in quantum field theory are predicated on acausal properties derived from general relativity. In this context, Blum et al. [6] stress the importance of the CPT theorem (quoting verbatim):
“CPT theorem is the statement that nothing would change—nobody would notice and the predictions of physics would not be altered—if we simultaneously replace particles by antiparticles and vice versa. Replace everything by its mirror image or, more exactly, exchange left and right, up and down, and front and back, and reverse the flow of time. We call this simultaneous transformation CPT, where C stands for Charge Conjugation (exchanging particles and antiparticles), P stands for parity (mirroring), and T stands for time reversal.”
4.2 Global pseudo-timelike loops
The existence of such pseudo-timelike curves locally near the hypersurface that loop back to themselves, gives rise to the question whether this type of curves also occur globally. We want to elucidate this question in the following. [^21] A spacetime is a Lorentzian manifold that models space and time in general relativity and physics. This is conventionally formalized by saying that a spacetime is a smooth connected time-orientable Lorentzian manifold $(M,g)$ with $\dim M=4$. But in what follows we want to study the $n$-dimensional ($n\geq 2$) case.
Definition 19 (Stably causal).
[31] A connected time-orientable Lorentzian manifold $(M,g)$ is said to be stably causal if there exists a nowhere-vanishing timelike vector field $V_{a}$ such that the Lorentzian metric on $M$ given by $g^{\prime}:=g_{ab}-V_{a}V_{b}$ admits no closed causal curves. In other words, if $(M,g)$ is stably causal then, for some timelike $V_{a}$, the metric $g^{\prime}:=g_{ab}-V_{a}V_{b}$ on $M$ is causal.
Remark 9.
A partial ordering << is defined in the set of all Lorentzian metrics $Lor(M)$ on $M$ in the following way: $g<g^{\prime}$ iff all causal vectors for $g$ are timelike for $g^{\prime}$. Then the metric $g_{\lambda}=g+\lambda(g^{\prime}-g)$, $\forall\>\lambda\in[0,1]$ is a Lorentzian metric on $M$, as well. Also, recall that $g<g^{\prime}$ means that the causal cones of $g$ are contained in the timelike cones of $g^{\prime}$. A connected time-orientable Lorentzian manifold $(M,g)$ is stably causal if there exists $g^{\prime}\in Lor(M)$, such that $g^{\prime}>g$, with $g^{\prime}$ causal.
Lemma 2.
[34] Stable causality is the necessary and sufficient condition for the existence of a smooth global time function, i.e. a differentiable map $T\colon M\rightarrow\mathbb{R}$ such that whenever $p<<q$ $\Longrightarrow T(p)<T(q)$.
Definition 20 (Globally hyperbolic).
[5, 24] A connected, time-orientable Lorentzian manifold $(M,g)$ is called globally hyperbolic if and only if it is diamond-compact and causal, i.e., $p\notin J^{+}(p)\>\forall p\in M$. [^22] Diamond-compact means $J(p,q):=J^{+}(p)\cap J^{-}(q)$ is compact for all $p,q\in M$. Note that $J(p,q)$ is possibly empty.
An equivalent condition for global hyperbolicity is as follows [15].
Definition 21.
A connected, time-orientable Lorentzian manifold $(M,g)$ is called globally hyperbolic if and only if $M$ contains a Cauchy surface. A Cauchy hypersurface in $M$ is a subset $S$ that is intersected exactly once by every inextensible timelike curve in MM. [^23] An inextensible curve is a general term that refers to a curve with no endpoints; it either extends infinitely or it closes in on itself to form a circle—a closed curve. Specifically, an inextensible timelike curve is a curve that remains timelike throughout its entire length and cannot be extended further within the spacetime. In mathematical terms, a map $\alpha\colon(a,b)\rightarrow M$ is an inextensible timelike curve in $(M,g)$ if $\alpha(t)$ does not approach a limit as $t$ increases to $b$ or decreases to $a$, and $\alpha(t)$ remains timelike for all $t\in(a,b)$. This distinguishes it from inextensible curves of other causal types, such as null or spacelike curves.
In 2003, Bernal and Sánchez [4] showed that any globally hyperbolic Lorentzian manifold $M$ admits a smooth spacelike Cauchy hypersurface $S$, and thus is diffeomorphic to the product of this Cauchy surface with $\mathbb{R}$, i.e. $M$ splits topologically as the product $\mathbb{R}\times S$. Specifically, a globally hyperbolic manifold is foliated by Cauchy surfaces.
Remark 10.
If $M$ is a smooth, connected time-orientable Lorentzian manifold with boundary, then we say it is globally hyperbolic if its interior is globally hyperbolic.
The next theorem (Global Loops Theorem) is partially based on the Local Loops Theorem 2 and can be considered a generalization to the global case.
Theorem (Global loops).
[…]
Proof.
[…]
Due to [4] we know that $M_{L}$ admits a splitting $M_{L}=(\mathbb{R}_{>0})_{t}\times S_{t}=\bigcup_{t\in\mathcal{\mathbb{R}}_{>0}}S_{t}$, such that the Lorentzian sector $M_{L}$ is decomposed into hypersurfaces (of dimension n1n-1), specified as the level surfaces $S_{t}=\mathcal{T}^{-1}(t)=\{p\in M_{L}\colon\mathcal{T}(p)=t\},t\in\mathbb{R}_{>0}$, of the real-valued smooth temporal function $\mathcal{T}\colon M_{L}\longrightarrow\mathbb{R}_{>0}$ whose gradient $\textrm{grad}\mathcal{T}$ is everywhere non-zero and, clearly, $d\mathcal{T}$ is an exact $1$-form. Within the neighborhood $U=\bigcup_{q\in\mathcal{H}}U(q)$ this foliation $\bigcup_{t\in\mathcal{\mathbb{R}}_{>0}}S_{t}$ can be chosen in such a way that it agrees with the natural foliation given by the absolute time function $\mathfrak{h}(t,\mathbf{\hat{x}}):=t$, see Remark 6 and Definition 16.242424Recall that a smooth function $T\colon M\longrightarrow\mathbb{R}$ on a connected time-orientable Lorentzian manifold $(M,g)$ is a global time function if $T$ is strictly increasing along each future-pointing non-spacelike curve. Moreover, a temporal function is a time function $T$ with a timelike gradient $\textrm{grad}T$ everywhere. Since $M_{L}$ is globally hyperbolic it admits a smooth global time function $T$ and consequently it admits [31] a temporal function $\mathcal{T}$. Hence, in the Lorentzian sector $M_{L}$ there exists a global temporal function $\mathcal{T}\colon M_{L}\longrightarrow\mathbb{R}_{>0}$, and $\textrm{grad}\mathcal{T}$ is orthogonal to each of the level surfaces $S_{t}=\mathcal{T}^{-1}(t)=\{p\in M_{L}\colon\mathcal{T}(p)=t\},t\in\mathbb{R}_{>0}$, of $\mathcal{T}$. Note that $\mathcal{T}=t$ is a scalar field on $M_{L}$, hence $\textrm{grad}\mathcal{T}=\textrm{grad}t=(dt)^{\#}$.
Moreover, the level surfaces $(S_{t})_{t\in\mathbb{R}}$ are Cauchy surfaces and, accordingly, each inextensible pseudo-timelike curve in $M_{L}$ can intersect each level set $S_{t}$ exactly once as $\mathcal{T}$ is strictly increasing along any future-pointing pseudo-timelike curve. [^25] Since $\mathcal{T}$ is regular the hypersurfaces $S_{t}$ never intersect, i.e. $S_{t}\cap S_{t^{\prime}}=\emptyset$ for $t\neq t^{\prime}$. Then, these level-sets $S_{t}$ are all space-like hypersurfaces which are orthogonal to a timelike and future-directed unit normal vector field nn. [^26] In other words, the unit vector $n$ is normal to each slice $S_{t}$, and $g$ restricted to $S_{t}$ is Riemannian.
For $\varepsilon$ sufficiently small, the level Cauchy surface
$S_{\varepsilon}=\mathcal{T}^{-1}(\varepsilon)=\{p\in M_{L}\colon\mathcal{T}(p)=\varepsilon\},\varepsilon\in\mathbb{R}_{>0}$
is contained in $U\cap M_{L}=\bigcup_{q\in\mathcal{H}}(U(q)\cap M_{L})$. [^27] This is true because all neighborhoods $U(q)$ with $q\in\mathcal{H}$ can be chosen such that the sets $U(q)$ have a compact closure. Thus, the $\overline{U(q)}$ are not “infinitely wide”, and there exists a strictly positive value $\varepsilon_{\max}$, such that for all $\varepsilon<\varepsilon_{\max}$, the level Cauchy surface $S_{\varepsilon}$ is contained in $U\cap M_{L}$.
Therefore, based on Theorem 2, for any $p=(\varepsilon,\mathbf{\hat{x}})\in S_{\varepsilon}\subseteq$$(U\cap M_{L})$ all past-directed and causal curves emanating from that point reach the hypersurface $\mathcal{H}$. The global hyperbolicity of $M_{L}$ implies that every non-spacelike curve in $M_{L}$ meets each $S_{t}$ once and exactly once since $S_{t}$ is a Cauchy surface. In particular, the spacelike hypersurface $S_{\varepsilon}$ is a Cauchy surface in the sense that for any $\bar{p}\in M_{L}$ in the future of $S_{\varepsilon}$, all past pseudo-timelike curves from $\bar{p}$ intersect $S_{\varepsilon}$. The same holds for all future directed pseudo-timelike curves from any point in $M_{L}$ in the past of $S_{\varepsilon}$.
Figure 8: In the future of the Cauchy surface $S_{\varepsilon}$, any point $\bar{p}\in M_{L}$ with $t>\varepsilon$ has all of its past-directed pseudo-timelike curves intersect $S_{\varepsilon}$. Similarly, all future-directed pseudo-timelike curves from a point $q\in\mathcal{H}$ intersect $S_{\varepsilon}$. This means there must be a point $\bar{q}\in\mathcal{H}$ from which $\bar{p}\in M_{L}$ can be reached by a future-directed pseudo-timelike curve.
Consequently, by virtue of Theorem 2 and the above argument, all past-directed pseudo-timelike curves emanating from any $\bar{p}\in M_{L}$ reach the hypersurface $\mathcal{H}$. Analogously we can conclude that any point $\bar{p}\in M_{L}$ can be reached by a future-directed pseudo-timelike curve starting at some suitable point in $\mathcal{H}$. Recall that, based on Remark 7, we also know that $\mathcal{I}^{+}(q)=\{p\in M:q\ll p\}=M$, that is, any point in $\color[rgb]{0,0,0}M=M_{R}\cup\mathcal{H}\cup M_{L}$} can be reached by a future-directed pseudo-timelike curve from $q\in\mathcal{H}$, see Figure 8.
We now obtain a loop with intersection point $p$ in $M_{L}$ if, for sufficiently small $\varepsilon$, we first prescribe the intersection point $p=(\varepsilon,\mathbf{\hat{x}})\in S_{\varepsilon}$. And then we connect the two points lying in $\mathcal{H}$ of the intersecting curve sections through an arbitrary curve segment in the Riemannian sector $M_{R}$ (through a suitable choice of the two curve segments, we can ensure that different points on $\mathcal{H}$ are obtained).
This reasoning also applies to points located on the hypersurface or within the Riemannian region. In these cases, the situation is as follows:
  1. 1.
    If the given point lies on the hypersurface, $p\in\mathcal{H}$, choose a timelike curve segment in $M_{L}$ connecting $p$ to $S_{\varepsilon}$ (with $\varepsilon$ sufficiently small). From there, continue along a second timelike curve segment back to another point on $\mathcal{H}$. Finally, connect both hypersurface points through the Riemannian sector. At $p$, the Riemannian curve segment must form a kink where it meets the timelike segment, and the two intersecting curve segments must be extended so that together they produce a self-intersection exactly at $p$.
  2. 2.
    If the given point lies in the Riemannian sector, $p\in M_{R}$, begin with a loop of the type constructed in the proof of Theorem 2. Modify the Riemannian portion of this loop so that it passes through $p$ and forms a self-intersection there. As in the previous case, one intersection point of the loop necessarily lies in $M_{L}$, while the second intersection is realized at $p$.
Remark 11.
Theorem 3 explicitly states that through every point in $M$, there always exists a pseudo-timelike loop. Consequently, this applies equally to points on the hypersurface as well as to those in the Riemannian region. In such cases, the construction necessarily produces a pseudo-timelike loop with two intersection points. Indeed, the point $p$ is required to serve as a point of self-intersection. Consequently, the pseudo-timelike loop must also admit an intersection in $M_{L}$, and at least one of the intersection points must lie in $M_{L}$ in order for the curve to qualify as a pseudo-timelike loop. The second intersection point is then determined, either as $p$ on $\mathcal{H}$ or as $p$ in $M_{R}$. [^28] The construction for points $p\in\mathcal{H}$ necessitates that the timelike curve segments from $M_{L}$ approach the “pinched” light cones near the hypersurface. Consequently, the curve segments that pass through $\mathcal{H}$ into the Riemannian sector $M_{R}$ and self-intersect at $p\in\mathcal{H}$ must meet with first-order contact. Topologically, a loop with two intersection points falls within the class of curves described in Definition 9.
Example 6.
The prototype of a spacetime $M$ with signature-type change is obtained by cutting an $S^{4}$ along its equator and joining it to the corresponding half of a de Sitter space. It is a well-known fact that the full de Sitter spacetime is globally hyperbolic [28], with the entire manifold possessing a Cauchy surface. When we restrict to half de Sitter space—by choosing an appropriate region bounded by a Cauchy surface—this region retains global hyperbolicity. This is because the Cauchy surface of the full de Sitter spacetime remains valid in the half-space, ensuring that every inextensible non-spacelike curve still intersects this surface exactly once. As a result, the Lorentzian sector, which corresponds to half de Sitter space, is also globally hyperbolic. Consequently, there are chronology-violating pseudo-timelike loops through each point in $M$.
Corollary 13.
[…]
5 Final thoughts
The intriguing facet of the potential existence of closed timelike curves within the framework of Einstein’s theory lies in the physical interpretation that CTCs, serving as the worldlines of observers, fundamentally permit an influence on the causal past. This can also be facilitated through a causal curve in the form of a loop, i.e., the curve intersects itself. In the case of a non-time-orientable manifold, there would then be the possibility that at the intersection, the two tangent vectors lie in different components of the light cone. Thus, the “time traveler” at the encounter with himself, which he experiences twice, may notice a reversal of the past and future time directions in his surroundings during the second occurrence, even including the behavior of his younger version. Regardless of whether this effect exists or not, during the second experience of the encounter, which he perceives as an encounter with a younger version of himself, the traveler can causally influence this younger version and its surroundings.
It is important to emphasize that the existence of pseudo-timelike loops near the hypersurface of signature change, as guaranteed by the Loop Theorems, does not imply a breakdown of physical causality. Locally in the Lorentzian region $M_{L}$, the directions of certain classes of curves are constrained, which can be interpreted physically as limitations on the propagation of signals or particle interactions along these curves. In contrast, in the Riemannian region $M_{R}$, two competing interpretations are possible regarding the physical meaning of curve segments:
  1. (1)
    Unrestricted propagation interpretation: While $M_{R}$ allows curves mathematically without directional constraints, one can treat them as a continuation of possible influence from $M_{L}$. Local causality in $M_{L}$ remains intact: in a sufficiently small neighborhood of a point $p$, no point $q$ can both influence and be influenced by $p$. In this reading, pseudo-timelike loops are mathematically extended curves that do not correspond to actual causal violation.
  2. (2)
    Causal barrier interpretation: Alternatively, $M_{R}$ can be viewed as a region where no physical propagation occurs. Curves there are spacelike in the Lorentzian sense and cannot carry influence. Consequently, any pseudo-timelike loop passing through $M_{R}$ cannot induce causality violations even in a global sense.
However, from the perspective of an observer in $M_{L}$ near the hypersurface $\mathcal{H}$, locally closed pseudo-timelike loops may appear as the creation of a particle–antiparticle pair at two distinct points $\hat{q},q\in\mathcal{H}$. One can visualize this as an object entering the Riemannian region, temporarily leaving the Lorentzian causal structure, and then re-emerging so as to propagate “forward in time” with respect to the time orientation in $M_{L}$. Since the mathematical definition of the loop itself does not privilege a direction of traversal, this picture is symmetric. This interpretation highlights that, while pseudo-timelike loops exist mathematically across the signature-changing hypersurface (see Subsection 4.1), they do not correspond to physically realizable violations of causality in $M_{L}$; instead, they provide a consistent physical picture in which $M_{R}$ either acts as a region in causal contact with $M_{L}$ or as a causal barrier.
Acknowledgments
The authors thank the reviewers for their insightful comments and critique, which have helped to refine and improve this article. NER is greatly indebted to Richard Schoen for generously welcoming her into his research group and to Alberto Cattaneo for affording her with creative independence throughout the duration of this research endeavor. Moreover, NER acknowledges partial support of the SNF Grant No. 200021-227719. This research was (partly) supported by the NCCR SwissMAP, funded by the Swiss National Science Foundation.
Data availability: No data was used for the research described in the article.
Conflict of interest
: The authors have no conflict of interest to declare that are relevant to the content of this article.